3(x-5)(x-7)=4(x+3)
Found 2 solutions by Boreal, solver91311:
Answer by Boreal(15235)

multiply out the left side and right
3(x^2-12x+35)=4x+12
3x^2-36x+105=4x+12
3x^2-40x+93=0
(x-3)(3x-31)=0
x=3, 31/3
first checks to 24=24
second checks to 160/3=160/3

Answer by solver91311(24713)


Rules for posting:

1. Ask an actual question. A question is a series of words that form a sentence where it is appropriate to append a question mark,(i.e. one of these things: "?") to the end of the sentence. Please look up the definition if you are still unclear on the concept. On this website, such question should relate to some area of mathematics.

2. If you post a problem for which you want a solution, make sure you make it clear what it is about the problem that is giving you difficulty.

3. Show all of the work, successful or not, that you have done or attempted in the solution of the problem.

4. If following these instructions is too much work for you, then helping you is too much work for us.


John

My calculator said it, I believe it, that settles it